Use the following collection of 30 test scores: 100, 65, 67, 61, 62, 70, 75, 73, 88, 77, 83, 79,95, 79, 8102, 86, 87, 87, 91, 87, 89, 92, 90, 99, 87, 72, 93, 79, 52

a. Construct a frequency distribution table using 6 classes: include classes, frequencies, class marks, class boundaries, relative frequencies, and relative percentages.

b. Find the mean to the nearest tenth and sample standard deviation to the nearest tenth.

c. Use the results from part b to determine the Z-score associated with a data value of 79.
Answer #1

Answer:-

Given That:-

Use the following collection of 30 test scores: 100, 65, 67, 61, 62, 70, 75, 73, 88, 77, 83, 79,95, 79, 8102, 86, 87, 87, 91, 87, 89, 92, 90, 99, 87, 72, 93, 79, 52

a. Construct a frequency distribution table using 6 classes: include classes, frequencies, class marks, class boundaries, relative frequencies, and relative percentages.

Class Class marks Frequencies Relative Frequencies Relative percentages
50-59 54.5 1 0.033 3.3
60-69 64.5 4 0.133 13.3
70-79 74.5 8 0.267 26.7
80-89 84.5 9 0.200 20
90-99 94.5 6 0.067 6.7
100-109 104.5 2

b. Find the mean to the nearest tenth and sample standard deviation to the nearest tenth.

Mean = 81.6

Stanadard deviation = 12.4

c. Use the results from part b to determine the Z-score associated with a data value of 79.

Z - score = (79 - 81.6)/12.4

= -0.21

(Mean and standard deviation are computed below)

The mean value of the data set is:

\mu=\frac{2447}{30}\approx 81.5667

Explanation:

The mean of a data set is the sum of the terms divided by the total number of terms. Using math notation we have:

Mean = Sum of terms/Number of terms

In this example:

Sum of terms = 100 + 65 + 67 + 61 + 62 + -----+ 52 = 2447

Number of terms = 30

Mean = Sum of terms/Number of terms

Mean = 2447/30

To find standard deviation we use the following formula

\sigma=\sqrt{\frac{\Sigma (x_i-\bar{X})^2}{n-1}}

we willl compute this formula in 4 steps.

Step 1:

Find the mean  (\bar{X})

In this example  \bar{X} = 81.5667. (use this calculator for step by step explanation on how to find mean)

Step 2:

Create the following table.

data data-mean (data-mean)2
100 18.4333 339.78654889
65 -16.5667 274.45554889
67 -14.5667 212.18874889
61 -20.5667 422.98914889
62 -19.5667 382.85574889
70 -11.5667 133.78854889
75 -6.5667 43.12154889
75 -6.5667 43.12154889
73 -8.5667 73.38834889
88 6.4333 41.38734889
77 -4.5667 20.85474889
83 1.4333 2.05434889
79 -2.5667 6.58794889
95 13.4333 180.45354889
79 -2.5667 6.58794889
80 -1.5667 2.45454889
102 20.4333 417.51974889
86 4.4333 19.65414889
87 5.4333 29.52074889
87 5.4333 29.52074889
91 9.4333 88.98714889
87 5.4333 29.52074889
89 7.4333 55.25394889
92 10.4333 108.85374889
90 8.4333 71.12054889
99 17.4333 303.91994889
87 5.4333 29.52074889
72 -9.5667 91.52174889
72 -9.5667 91.52174889
93 11.4333 130.72034889
79 -2.5667 6.58794889
52 -29.5667 847.18974889

Step 3:

Find the sum of numbers in the last column to get.

\sum (x_i-\bar{X})^2=4429.3667

step 4:

Calculate  \sigma using the baove formula.

\sigma=\sqrt{\frac{\Sigma (x_i-\bar{X})^2}{n-1}}

\sigma=\sqrt{\frac{4429.3667}{30-1}}

\sigma\approx 12.3587

Thank you for your supporting. Please upvote my answer...